OpenStudy (anonymous):

A bit of explanation here, s is a distance with unit [L] and theta is an angle in radians, I need to find c in terms of [L] and [T] (T being a unit of time) for this equation "theta=s/c" if you can explain how to remove theta without a sin or cos popping up on the other side i would greatly appreciate it.

OpenStudy (e.mccormick):

so \(\dfrac{s}{c}\) is measured in radians. Hmmm. This sounds like some sort of arc happening over time. Have you looked at the arc measure formulas?

OpenStudy (e.mccormick):

\(\theta = \dfrac{ArcLen}{radius}\)

OpenStudy (anonymous):

Well none of my answers have trig functions in them, one answer is [c]=1/[L]. Answers like that i just don't see how you cam remove the theta without using some kind of trig function.

OpenStudy (anonymous):

son of a gun your a genius, I see what you mean now!

OpenStudy (e.mccormick):

Nah, I just thought about what other formulas used theta in radians but not the trig ratios. At least it gives you a starting point. Then you get to play algebra games to solve the rest.

OpenStudy (anonymous):

Yeah, thanks and i got the answer thanks a ton.

OpenStudy (e.mccormick):

np. have fun!